OpenStudy (anonymous):

Find the equation (in terms of x ) of the line through the points (-3,5) and (4,1)

OpenStudy (anonymous):

the equation of the line that passes through two points (x1,y1) and (x2,y2) is\[\frac{y - y_1}{y_2 - y_1} = \frac{x - x_1}{x_2 - x_1}\]

OpenStudy (anonymous):

looks like you got it good luck

OpenStudy (anonymous):

so (y-5)/-4=(x+3)/7?

OpenStudy (anonymous):